OpenStudy (anonymous):

just a quick question: Consider the solid that lies above the square (in the xy-plane) R=[0,2]×[0,2], and below the elliptic paraboloid z=49−x2−2y2. Using iterated integrals, compute the exact value of the volume. the R in this...is that my bounds?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

R is the shadow of surface `z = f(x,y)` in xy plane

ganeshie8 (ganeshie8):

\[\large V = \int\limits_{0}^2\int\limits_0^2 49-x^2-2y^2~dx~dy\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\int\limits_{0}^{2}49-x^2-2y^2\]

OpenStudy (anonymous):

forgot the dxdy. but yeah. that is an easy problem. just wasn't sure what the bounds were

OpenStudy (anonymous):

thanks!